OpenStudy (anonymous):

The price of a technology stock was $9.64 yesterday. Today, the price rose to $9.77 . Find the percentage increase. Round your answer to the nearest tenth of a percent.

OpenStudy (anonymous):

$9.77 -$9.64 =0.13\[\frac{0.13}{9.64}=\frac{x}{100}\] cross n time (0.13*100)=9.64x 13/9.64 = x 1.35=x If I deserve medal then give me one.